OpenStudy (anonymous):

lim ( (1/x+2)^2-1/4 )/x , evaluate the limit or x->0 state that it Does Not Exist

OpenStudy (lgbasallote):

\[\huge\lim _{x \rightarrow 0} \frac{(\frac 1x + 2)^2 - \frac 14}x\] is that your question?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

do you see that your numerator is a difference of two squares?

OpenStudy (anonymous):

I don't know if my first step should be to write \[(1x^{-1}-\frac{ 1 }{ 4 })\]